OpenStudy (anonymous):

the product of four and a number is 25. what is the number? please show work

OpenStudy (anonymous):

the number is 6.25

OpenStudy (anonymous):

6.25x4=25

OpenStudy (anonymous):

4x=25 (4x)/4=25/4 x=25/4